OpenStudy (anonymous):

A 1500 kg car is pulled by a force of 4,000 N. What is the acceleration of the car?

OpenStudy (anonymous):

F=ma F=4000N m=1500 a=? a=F/m= 4000N/1500kg = 2.67m/s^2